OpenStudy (anonymous):

Can someone tell me if this is right please. Find the area of the indicated region. Enclosed by y = 2^x, y = x + 2, x = −2, and x = 2 ...

OpenStudy (anonymous):

\[\int\limits_{-2}^{-1.69}2^x-x+2\] \[\int\limits_{-1.69}^{2} x+2-2^x\]

OpenStudy (anonymous):

if you say "indicated region", then a graph is expected.

OpenStudy (anonymous):

 graph looks like this

OpenStudy (anonymous):

ooh. its \(y=2^x\)

OpenStudy (anonymous):